how to solve this system of equation graphically,
x+y=-4
x-y=6

Answer by checkley77(12844)

x+y=-4
x-y=6 add them.
------------
2x=2
x=2/2
x=1 ans.
1+y=-4
y=-4-1
y=-5 ans.
Proof:
1-(-5)=6
1+5=6
6=6